Two thirds Kay's age increases by 12 is equal to twice her age decreased by 3. How old is Kay?
Answer by stanbon(75887)

Two thirds Kay's age increased by 12 is equal to twice her age decreased by 3. How old is Kay?
-------
Equation:
(2/3)(k + 12) = 2(k-3)
----
(2/3)k + 8 = 2k-6
(4/3)k = 14
k = (3/4)14
k = 42/4
k = 10 1/2 years
------
Cheers,
Stan H.
---------------